OpenStudy (anonymous):

2 over 3 = 4 over x?

OpenStudy (anonymous):

Cross-multiply for starters:\[\Large \frac{a}{b} = \frac{c}{d} \longrightarrow ad=bc\]

OpenStudy (lgbasallote):

applying that to your problem.. \[\frac{2}{3} = \frac{4}{x}\] \[\Rightarrow 2x = 4 \times 3\] \[\Rightarrow 2x = 12\] can you solve for x now?

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

try dividing BOTH sides by 2 so that 2x will just become x

OpenStudy (lgbasallote):

what will happen to 12 if you divide it by 2?

OpenStudy (theviper):

Do ur problem by cross multiplication.:)

OpenStudy (lgbasallote):

@krizzypoo12 are you there?

OpenStudy (theviper):

\[x=\frac{12}{2}=6\]

OpenStudy (theviper):

Got it???? I think now u can get it:)

OpenStudy (anonymous):

yeah im here